Solve the following equation and check your answer:
3(2x-2)-4(x+6) = 3x+8+x
6x -6 -4x -24 = 4x+8
2x-30 = 4x+8
-4x-30 = -4x+30
= 2x = -38
-2 2
x= -19
Is this the correct way of solving the equation? How do I check my answer?

Answer by jim_thompson5910(35256)

Your answer and steps are correct. Congrats on that.

Here's the check.

Check:

3(2x-2)-4(x+6) = 3x+8+x

3(2(-19)-2)-4(-19+6) = 3(-19)+8+(-19) ... Plug in x = -19

3(-38-2)-4(-19+6) = -57+8-19

3(-40)-4(-13) = -57+8-19

-120+52 = -57+8-19

-68 = -68

Since that last equation is true, this verifies the solution.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------